OpenStudy (anonymous):

Find the Surface are of a sphere whose great circle have a circumference of 8 pie (The symbol )

OpenStudy (anonymous):

this is the actual Question, Find the Surface area of a sphere whose great circle have a circumference of 8 pie (symbol: 3.14)

OpenStudy (anonymous):

The circumference \[2 \pi radius= 8 \pi\] (dividing by 2pi) so the radius=4 The surface area of a sphere= \[4 \pi r^2= 4 \pi (4)^2=64 \pi\]

OpenStudy (anonymous):

thanks! :D

OpenStudy (anonymous):

you're welcome